Aluminium Nitrate reacts with Sod. Hydroxide to form Aluminium Hydroxide. What mass of Aluminium hydroxide will ppt.
When 35.5 ml of 0.145 mol/L aq. Aluminium Nitrate is mixed with 44.2mlof 0.215 mol/L Aq. Sod. Hydoxide? Kindly provide solution)
1
Expert's answer
2020-12-06T13:36:21-0500

In this reaction, NaOH is the limiting reagent

Mole of NaOH = 44.2 × 0.01L × 0.215M = 0.009503 mol

Al(NO3)3 + 3NaOH = Al(OH)3 + 3NaNO3

3 mole NaOH = 1 mole Al(OH)3 = 78g Al(OH)3

1 mole NaOH = 78/3 g Al(OH)3

0.009503 mole of NaOH = 78 × 0.009503/ 3 g Al(OH)3

= 0.25g Al(OH)3